### v1

@@ Abstract
In addition to the oval-shaped body, the Late-stage sea star larvae also have protruding arms which act as localized sources of vorticity generation. Hence, under strong confinement, these larvae generate three pairs of vortices (Pairs 1,2,3) (Fig. 2 d, left panel). The theoretical model considering three pairs of Stokeslets placed exactly adjacent to the arms corresponding to the physical locations, once again shows excellent agreement with the experiments (Fig. 2d, right panel).
@@
In the Sea urchin larvae, in addition to the tips of the extended arms, a hood region just below the arms also act as localized sources of vorticity generation. These morphological complexities enabled these larvae to generate three pairs of vortices (Pairs 1,2,3) under strong confinement (Fig. 2f left panel). To develop these three pairs of vortices in theory, a 2 pair Stokeslet model is used. In this context, one pair of Stokeslets are introduced at the back of the sea urchin larvae that accounted for the hood. Next, at the tip of each arm, one Stokeslet was introduced. Overall, this two pair Stokeslet model is able to capture the Sea urchin larval experimental flow-fields very well ((Fig. 2f left panel)).
@@
Overall, the theoretical model using superposition of Stokeslets qualitatively captured the experimental results very well in all the three morphologically complex larvae in different confinement conditions, ranging from weak to high squeeze confinement. Further, the theoretical model was compared quantitatively with the experiments by calculating the magnitude of spatial velocity decay over the vortices (Fig. 2 g-j). Under weak confinement conditions (), since all the three larvae generate only one pair of vortices (Pair 1), we compared their velocity magnitude decay through the two vortex centers (averaged over Pair 1) from both the experiments and theoretical model in Fig. 2g. For the three types of larvae, the velocity at the larval body surface starts in the range 50-150 m and decays to a minimum (sharp dip representing the vortex center), and eventually decays to about half of its magnitude (30-70 m ) over a length scale of about 1 (Fig. 2g). For the Early-stage sea star, the theoretical model matches the experiments almost exactly, and for the Late-stage sea star and Sea urchin, the model is unable to match the experiments quantitatively, but still follows the overall trends. The velocity decays for all the types of larvae follow slopes in the range between and . The initial velocity decay is closer to , and after crossing the velocity dip at the vortex center, the decay becomes less steep with a slope closer to (Fig. 2g).
@@
For strong confinement (), the spatial velocity magnitude decays between the experiments and the theoretical model for the three types of larvae were compared for each vortex pair (Pair 1,2,3) separately (Fig. 2 h-j). The theoretical model has the best match for the Early-stage sea star, followed by the Sea urchin, and then the Late-stage sea star larvae. The experimental velocity decays from a range of 8-35 m at the surface to about half the value (3-15 m ) for the Early-stage sea star and Sea urchin larvae (Fig. 2 h,j), but this trend does not hold for all vortex pairs in the Late-stage sea star larvae (Fig. 2 i). In all the three types of larvae, the theoretical model shows a steeper initial velocity decay of compared to the experimental data, but the decay slope of towards the end matches the experimental data (Fig. 2 h-j). In all the larvae, the model predicts a sharp dip in velocity at the vortex center, but only the Early-stage sea star experiments show this sharp dip (Fig. 2 h).
@@
Our theoretical model captures the experimental velocity decay trends reasonably well, but not the magnitudes. For the three types of larvae, in both the weak and strong confinement conditions, although the theoretical model velocity decay slopes range between and , the experimental data more closely follow the decay (Fig. 2 g-j). In comparison, the velocity decays in the low Reynolds numbers flows in Chlamydomonas reinhardtii have been found to decay as [26], which is also the expected velocity decay for a point source confined between two parallel no-slip walls [6, 26]. The less steep velocity decays observed in the present work, in the range between and , can most likely be attributed to inertial effects. Future theoretical work in this regime of low-to-intermediate Reynolds numbers (0.1–0.9) is necessary to deepen our understanding of the physical mechanisms underlying the velocity decays.
@@
The theoretical model described above using superposition of Stokeslets matched the experimental confinement-induced flows (Fig. 2) and revealed two key features of confinement-induced flows in the morphologically complex larvae of Sea stars and Sea urchin. First, under strong confinement, the theoretical model utilized a unit Stokeslet to represent a localized source of vorticity generation in the experiments, often corresponding to a morphological location on the larval body such as the convex curvature change in the Early-stage sea star larva, the protruding arms in Late-stage sea star larva, or the tip of extended arms in Sea urchin larvae (Fig. 2 b,d,f). Second, a quantitative comparison of the experimental confinement-induced velocity decays for all the types of larvae (Fig. 2 g-j) revealed that the velocity magnitudes under weak confinement are about five times higher than the velocity magnitudes under strong confinement. This result suggests that frictional effects from the bottom and top walls decrease the velocity magnitude as the walls come closer due to increase in confinement. Putting together these two key features, we present a general framework to explain the increase in number of vortices due to increase in confinement.
@@
In the example of the Late-stage stage sea star larva, the body morphology consists of an extended oval/ellipsoidal shape with protruding arms. This Late-stage sea star larva generates two vortices under weak confinement (Fig. 1c, Fig. 2c), four vortices under moderate confinement (Extended Fig. 1), and six vortices under strong confinement (Fig. 1d, Fig. 2d). In Fig. 3 a-c, the vorticity field generated by the Late-stage sea star larva is reexamined for weak, moderate and strong confinements. The three source pairs (Pair 1,2,3) of local vorticity generation are visible in all the three confinement conditions. In weak confinement, the Pair 1 vortices dominate the flow-field with the largest vortex size (area) (Fig. 3d) (Extended Data Fig. 5 d) (Methods). However, the local sources of vorticity generation (Pairs 2 and 3) corresponding to the locations of the protruding arms are visible as small patches of vorticity (Fig. 3a,d). In moderate confinement, the frictional effects from the top and bottom walls are increased, which decreases the size of the dominant Pair 1 vortices (Fig. 3b,d), and in turn provides an opportunity for Pair 2 and 3 vortices to increase in size (Fig. 3d). However, only Pair 2 vortices are able to develop properly given that these are located far away from Pair 1. Hence, Pairs 1 and 2 dominate the flow field and suppress the formation of Pair 3 vortices (Fig. 3b). Under strong confinement, the wall frictional effects are further increased, and this further decreases the size of Pair 1 vortices (Fig. 3d), and enables the generation of Pair 3 vortices. Hence, the flow-field is dominated by all the three vortex Pairs (1,2,3), giving rise to a total of six vortices. The vortex size results for Early-stage sea star and Sea urchin larvae under weak, moderate, and strong confinement also show the same trend observed in Late-stage sea star larvae (Fig. 3e) (Extended Data Fig. 4)(Extended Data Fig. 5). Hence, the confinement-induced flow-fields show similar trends for the larvae of very different forms and local morphologies. This suggests that the phenomenon of increase in number of vortices due to increase in confinement, only depends on the number of sources of local vorticity generation that can either get suppressed or amplified depending on whether the confinement is weak or strong. Importantly, this phenomenon is independent of details of the overall global larval form but only dependent on the local sources of vorticity generation (local morphological features such as tips of arms, sharp convex regions, etc.).
@@
In summary, we investigated the flow-fields induced by squeeze-confinement at low-to-intermediate Reynolds numbers (1) in the distinct larval forms of Early-stage and Late-stage sea stars, and Sea urchins. In all three cases, the larvae generated two vortices under weak confinement, but under strong confinement they developed multiple vortices. A low Reynolds number-based theoretical model utilizing superposition of multiple Stokeslets showed good qualitative agreement with the experiments. Based on the experiments and theory, we constructed a framework to explain the universality of vortex dynamics under confinement, highlighting the importance of morphological features inducing local vorticity generation.
@@
Hydrodynamics resulting from microorganism-surface interactions are intriguing and often give rise to non-intuitive physical phenomena. While the present work elucidates the fluid dynamics associated with larval squeeze-confinement, alternative confinement methods—such as tethering [21, 24, 25] or strong adhesive attachment [35]—could significantly alter the boundary conditions and lead to novel fluid dynamic effects. Also, the larval flows in the current work are in the regime of low-to-intermediate Reynolds numbers (0.1-0.9), so even though viscous effects are more important, inertial effects also need to considered. Although our theoretical model based on low Reynolds numbers [7, 6] performs well qualitatively, it does not work quantitatively since the model does not account for inertial effects. Hence, there is a need for the development of a new class of theoretical microswimmer models in this low-to-intermediate Reynolds numbers regime.
@@
Confinement-induced fluid dynamical effects on microorganisms can also have important biological consequences [5, 17]. Although it is unlikely that echinoderm larvae encounter squeeze-confinement conditions in the ocean, they are likely to interact with surfaces, and these hydrodynamic interactions can influence their swimming and feeding behaviors [5], which will be the focus of our future studies. Thus, while our fluid dynamics results are generally relevant to microswimmer-surface interactions, they are particularly applicable to quantitative laboratory-based studies involving microscopic confinement of a wide range of ciliated and motile marine microorganisms with varied forms and morphologies [27, 17], including collective phenomena that emerge when large numbers of larvae interact in close proximity [36, 37]. Beyond this, there are several potential applications in both biology and engineering. For instance, our findings could aid in the development of quantitative assays for studying larval morphological changes in response to chemicals [38], as well as ecological assays to assess microplastics accumulation in marine zooplankton [39, 40]. In engineering applications, these insights from ciliary flows could contribute to the design of microfluidic devices like lab-on-a-chip systems[41], artificial water purification devices [42, 43], and morphologically optimized micro-robots for targeted drug delivery [44], environmental monitoring [45], and beyond.
@@ Methods
Laboratory cultures of marine larvae and algae
Adult sea stars (bat stars, Patiria miniata) were collected and shipped to Miami from the California coast (Marinus Scientific, CA), and males and females were kept separately in two 20-gallon glass aquarium tanks. The aquariums were assembled from commercially available standard components (Petco). Each aquarium tank was filled with seawater (Imagitarium pacific ocean water, Petco) and maintained at a temperature of C using an external chiller unit. The tank seawater was continuously circulated using a pump and filtered, and regular water changes ( replacement) were carried out monthly. The sea stars were fed cut pieces of supermarket-purchased shrimp once a month.
The larvae were fed with marine algae, both red algae (Rhodomonas lens, Rhodomonas salina) and green algae (Dunaliella tertiolecta), which were obtained from the National Center for Marine Algae and Microbiota (NCMA, Bigelow, ME). The algae were cultured in erlenmeyer flasks containing filtered seawater and nutrients (Micro Algae Grow, Florida Aqua Farms) in 1:100 dilution and kept in an incubator maintained at a temperature of C, under 12 hour dark/light cycles using LED lighting. The algae flasks were shaken at least once daily to prevent settling at the bottom, and the algae cultures were split weekly into 1:4 volumes of fresh seawater and nutrients.
Spawning of marine larvae
In-vitro fertilization of the sea star (Patiria miniata) gametes were carried out in glass dishes. The eggs collected from the female sea stars were kept in 100 M 1-methyladenine (1-MA) to mature eggs for about 30-60 minutes. Then freshly collected sperm from the males were added (1:1000 dilution) and gently mixed in the seawater using a plastic pipette. After successful fertilization, the glass dishes containing growing embryos were placed in incubators maintained at a temperature of C. The red and green marine algae R. lens, R. salina, D. tertiolecta) were provided as a food source for growing larvae, and regular water changes were carried out weekly. The sea urchin larvae (Lytechinus variegatus) were obtained from the animal culture facility at The National Resource for Aplysia on the campus of University of Miami’s Rosenstiel School of Marine, Atmospheric, and Earth Science.
@@
A larva was confined or squeezed gently between a bottom glass slide and a top cover slip with different heights (z-direction) using a spacer tape, and fluid flow-fields were quantified in a quasi-2D x-y plane. Glass slides were first prepared by attaching a pair of long rectangular strips of double-sided tape of different thicknesses 50/100/150 m (Nitto Inc.) on the slide edges to serve as spacers for a required boundary height (H). Using this method, the boundary height (H) was varied systematically up to 900 m, with minimum increments of 50 m.
A small droplet (L) of seawater containing the larva of interest was transferred using a micropipette from the culture dish into the center of the prepared glass slide with the spacers at the edges. A smaller droplet ( L) containing m-sized tracer polystyrene microspheres (Polysciences Inc, with 1:100 dilution in seawater) was added to the droplet with the larva and gently mixed.
A cover slip was introduced slowly onto the glass slide containing the larva and microspheres, and attached on the spacer tapes at the edges. Next, live-imaging experiments were carried out by mounting this glass slide on an upright microscope (Zeiss Axio Imager M2). The microsphere movements were imaged using darkfield time-lapse imaging using a 10X objective to capture a 1.5 mm 1.5 mm field of view. The time-lapse images were acquired at 10 fps for a duration of 30 s using a high-speed camera (Hamamatsu ORCA-Fusion Gen-III sCMOS) attached to the Zeiss microscope.
@@
The larval flow-fields were qualitatively visualized using the Flowtrace technique [32]. The flowtrace analysis was run using the image-J plugin for 30 s durations over a time-averaging window of 3 s. Flowtrace visualizations were carried out on all the experimental datasets, and these videos were used to count the number of vortices (number of closed circular pathlines) for each case of confinement. The number of vortices was noted for the entire 30 s time sequence of a single dataset, and then an average was taken over at least N = 3 samples to represent the average number of vortices for each case of confinement and larva. The total number of datasets analyzed for Early-stage sea stars, Late-stage sea stars and Sea urchin larvae are N = 55, 44, and 26 respectively (Table 1, Supplementary Information).
The Particle Image Velocimetry (PIV) technique was adopted for quantification of the experimental larval flows. The time-lapse images were post-processed using PIVlab in MATLAB [33]. A mask was drawn in PIVlab over the larva to avoid any flow quantification noisy artifacts inside the larval body. The time-lapse images were first pre-processed to remove background noise using high-pass filters. Then the images were subjected to the PIV analysis where the time-varying velocity vector fields are computed. These velocity fields were further processed to smoothen and interpolate the vector fields as necessary. Next, the flow parameters such as vorticity or velocity magnitude were plotted along with the velocity vector fields. The averaged mean of parameters such as vorticity or velocity magnitude were calculated over a time window of 3 s in order to obtain smooth fluid flow visualizations and quantification. The distance from the center of vortices to the larval body surface was measured in N = 3 larval datasets and averaged to obtain the average distance measurements.
@@
In order to theoretically model fluid flows generated by the ciliated marine larvae under squeeze-confinement, we used an approach utilizing superposition of Stokeslets [7, 6]. In the experiments, the ciliated larvae were confined between two plates and the flows were measured at the midplane, i.e. z = H/2. Recently, flow fields generated by a Chlamydomonas confined between two parallel plates was studied using both experiments and theory [7]. Inspired by this work, we extended the theoretical model and applied to our experiments in ciliated marine larvae. The theoretical model starts with the incompressible 3D Stokes equation (as in Ref [7]),
@@
where is the point source of force generation in the z = H/2 plane. The solution to the above equation (using Fourier analysis) is obtained as:
The velocity fields are numerically solved over a 30 x 30 x-y grid in MATLAB to obtain the theoretical flow-fields. In Ref [7], the Chlamydomonas has two flagella that beat to generate vortices, however, here the marine larvae generate fluid vortices using a large number of cilia located on their ciliary bands. Although the methods of inducing fluid flow are different in the two studies, the consequence, i.e. vortex generation, is the same in both cases. Here, the focus is on modeling the overall flow-fields induced by confinement rather than the local mechanism of vorticity generation.
@@
Videos 1 and 2.
PDF file with supporting text, Figures S1-S4, Tables SX-SX, and Video Legends.

### v2

@@ Abstract
In addition to the oval-shaped body, the late stage sea star larvae also have protrusions which act as localized sources of vorticity generation. Hence, under strong confinement, these larvae generate three pairs of vortices (Pairs 1, 2, 3) (Fig. 2d, left panel). The theoretical model considering three pairs of Stokeslets placed exactly adjacent to the these protrusions, once again shows excellent agreement with the experiments (Fig. 2d, right panel).
@@
In the sea urchin larvae, the tips of the extended arms and a hood region just below the arms also act as localized sources of vorticity generation. These local morphological complexities enabled larvae to generate three pairs of vortices (Pairs 1, 2, 3) under strong confinement (Fig. 2f left panel). To develop these three pairs of vortices in theory, a 2 pair Stokeslet model is used. In this context, one pair of Stokeslets are introduced at the back of the sea urchin larvae that accounted for the hood. Next, at the tip of each arm, one Stokeslet was introduced. Overall, this two pair Stokeslet model is able to capture the sea urchin larval experimental flow fields very well (Fig. 2f left panel).
@@
Overall, the theoretical model using superposition of Stokeslets qualitatively captured the experimental results very well in all three morphologically complex larvae under different confinement conditions, ranging from weak to high squeeze confinement. Further, the theoretical model was compared quantitatively with the experiments by calculating the magnitude of spatial velocity decay over the vortices (Fig. 2g-j). Under weak confinement conditions () , since all the three larvae generate only one pair of vortices (Pair 1), we compared their velocity magnitude decay through the two vortex centers (averaged over Pair 1) from both the experiments and theoretical model in Fig. 2g. For the three types of larvae, the velocity at the larval body surface starts in the range 50-150 m and decays to a minimum (sharp dip representing the vortex center), and eventually decays to about half of its magnitude (30-70 m ) over a length scale of about m (Fig. 2g). For the early stage sea star, the theoretical model matches the experiments almost exactly, and for the late stage sea star and sea urchin, the model is unable to match the experiments quantitatively, but still follows the overall trends. The velocity decays for all the types of larvae follow slopes in the range between and . The initial velocity decay is closer to , and after crossing the velocity dip at the vortex center, the decay becomes less steep with a slope closer to (Fig. 2g).
@@
For strong confinement (), the spatial velocity magnitude decays between the experiments and the theoretical model for the three types of larvae were compared for each vortex pair (Pair 1, 2, 3) separately (Fig. 2h-j). The theoretical model has the best match for the early stage sea star, followed by the sea urchin, and then the late stage sea star larvae. The experimental velocity decays from a range of 8-35 m at the surface to about half the value (3-15 m ) for the early stage sea star and sea urchin larvae (Fig. 2h, j), but this trend does not hold for all vortex pairs in the late stage sea star larvae (Fig. 2i). In all the three types of larvae, the theoretical model shows a steeper initial velocity decay of compared to the experimental data, but the decay slope of towards the end matches the experimental data (Fig. 2h-j). In all larvae, the model predicts a sharp dip in velocity at the vortex center, but only the early stage sea star experiments show this sharp dip (Fig. 2h).
@@
Our theoretical model captures the experimental velocity decay trends reasonably well, but not the magnitudes. For the three types of larvae, in both the weak and strong confinement conditions, although the theoretical model velocity decay slopes range between and , the experimental data more closely follow the decay (Fig. 2g-j). In comparison, the velocity decays in the low Reynolds numbers flows in Chlamydomonas reinhardtii have been found to decay as [26], which is also the expected velocity decay for a point source confined between two parallel no-slip walls [6, 26]. The less steep velocity decays observed in the present work, in the range between and , can most likely be attributed to inertial effects since . Future theoretical work in this regime of low-to-intermediate Reynolds numbers (0.1–0.9) is necessary to deepen our understanding of the physical mechanisms underlying the velocity decays.
@@
The theoretical model described above using superposition of Stokeslets matched the experimental confinement-induced flows and revealed two key features of confinement-induced flows in the morphologically complex larvae of sea stars and sea urchins. First, under strong confinement (), the theoretical model utilized a unit Stokeslet to represent a localized source of vorticity generation in the experiments, corresponding to a morphological location on the larval body such as the convex curvature change in the early stage sea star larvae, the protrusions in late stage sea star larvae, or the tip of extended arms in sea urchin larvae (Fig. 2b, d, f). Second, a quantitative comparison of the experimental confinement-induced velocity decays for all the types of larvae (Fig. 2g-j) revealed that the velocity magnitudes under weak confinement () are about five times higher than the velocity magnitudes under strong confinement. This result suggests that frictional effects from the bottom and top walls decrease the velocity magnitude as the walls come closer due to the increase in confinement. Putting together these two key features, we present a general framework to explain the increase in number of vortices due to the increase in confinement.
@@
In the example of the late stage stage sea star larva, the body morphology consists of an extended oval/ellipsoidal shape with protruding arms. This late stage sea star larva generates two vortices under weak confinement (, Fig. 1c, Fig. 2c), four vortices under moderate confinement (, Extended Fig. 1), and six vortices under strong confinement (, Fig. 1d, Fig. 2d). In Fig. 3a-c, the vorticity field generated by the late stage sea star larva is reexamined for weak, moderate and strong confinements. The three source pairs (Pair 1, 2, 3) of local vorticity generation are visible in all the three confinement conditions. In weak confinement, the Pair 1 vortices dominate the flow-field with the largest vortex size (area) (Fig. 3d) (Extended Data Fig. 6d) (Methods). However, the local sources of vorticity generation corresponding to the locations of the protruding arms (Pairs 2 and 3) are visible as small patches of vorticity (Fig. 3a, d). In moderate confinement, the frictional effects from the top and bottom walls are increased, which decreases the size of the dominant Pair 1 vortices (Fig. 3b, d), and in turn provides an opportunity for Pair 2 and 3 vortices to increase in size (Fig. 3d). However, only Pair 2 vortices develop properly given that these are located far away from Pair 1. Hence, Pairs 1 and 2 dominate the flow field and suppress the formation of Pair 3 vortices (Fig. 3b). Under strong confinement, the wall frictional effects are further increased, and this further decreases the size of Pair 1 vortices (Fig. 3d), and enables the generation of Pair 3 vortices. Hence, the flow field is dominated by all the three vortex Pairs (1, 2, 3), giving rise to a total of six vortices. This relationship between vortex size, number, and degree of confinement – two, large vortices under low confinement and many, small vortices under high confinement – is present in all larvae tested, suggesting similar trends across different forms and local morphologies. (Fig. 3e) (Extended Data Fig. 4)(Extended Data Fig. 6). It follows from this that the number of observable vortices depends on the number of sources of local vorticity generation that are either suppressed or amplified depending on if the confinement is weak or strong, independent of the overall larval body plan and only dependent on the local sources of vorticity generation (local morphological features such as tips of arms, sharp convex regions, etc.).
@@
In summary, we investigated the flow fields induced by squeeze-confinement at low-to-intermediate Reynolds numbers ( 1) in the distinct larval forms of early stage and late stage sea star and sea urchin larvae. In all three cases, larvae generated two vortices under weak confinement, but under strong confinement they developed multiple vortices. A low Reynolds number-based theoretical model utilizing superposition of multiple Stokeslets showed good qualitative agreement with the experiments. Based on the experiments and theory, we constructed a framework to explain the universality of vortex dynamics under confinement, highlighting the importance of local morphological features inducing local vorticity generation.
@@
Hydrodynamics resulting from microorganism-surface interactions are intriguing and often give rise to non-intuitive physical phenomena. While the present work elucidates the fluid dynamics associated with larval squeeze-confinement, alternative confinement methods—such as tethering [21, 24, 25] or strong adhesive attachment [35]—could drastically alter the boundary conditions and lead to novel fluid dynamic effects. Also, the larval flows in the current work are in the regime of low-to-intermediate Reynolds numbers (0.1-0.9), so even though viscous effects are more important, inertial effects also need to considered. Although our theoretical model based on low Reynolds numbers [7, 6] performs well qualitatively, it does not work quantitatively since the model does not account for inertial effects. Hence, there is a need for the development of a new class of theoretical microswimmer models in this low-to-intermediate Reynolds numbers regime.
@@
Confinement-induced fluid dynamical effects on microorganisms can also have important biological consequences [5, 17]. Although it is unlikely that echinoderm larvae encounter squeeze-confinement conditions in the ocean, they are likely to interact with surfaces and these hydrodynamic interactions can influence their swimming and feeding behaviors [5], which will be the focus of our future studies. Thus, while our fluid dynamics results are generally relevant to microswimmer-surface interactions, they are particularly applicable to quantitative laboratory-based studies involving microscopic confinement of a wide range of ciliated and motile marine microorganisms with varied forms and morphologies [27, 17], including collective phenomena that emerge when large numbers of larvae interact in close proximity [36, 37]. Beyond this, there are several potential applications in both biology and engineering. For instance, our findings could aid in the development of quantitative assays for studying larval morphological changes in response to chemicals [38], as well as ecological assays to assess microplastics accumulation in marine zooplankton [39, 40]. In engineering applications, these insights from ciliary flows could contribute to the design of microfluidic devices like lab-on-a-chip systems[41], artificial water purification devices [42, 43], and morphologically optimized micro-robots for targeted drug delivery [44], environmental monitoring [45], and beyond.
@@ Methods
Laboratory cultures, larval spawning, and maintenance
Adult sea stars (Patiria miniata) were collected and shipped to the University of Miami (Miami, FL, USA) from the California coast (Marinus Scientific, CA). Ten male and ten female sea stars were kept separately in two 20 gal glass aquarium tanks. Aquariums were filled with seawater (Imagitarium pacific ocean water, Petco) and maintained at a temperature of C using an external chiller unit. The tank water was continuously circulated using a pump and filtered. Tanks underwent monthly water changes. The sea stars were hand-fed cut pieces of supermarket-purchased shrimp once a month, with each adult animal consuming approximately 1/3 of a medium-sized shrimp.
In-vitro fertilization of sea star gametes were carried out in large glass ovenware dishes (3 qt) in seawater using standard protocols [46]. The gametes were collected by making a 2 cm lateral incision on the ventral side of the sea star arm parallel to the lateral canal. The eggs collected from female sea stars were kept in 100 M 1-methyladenine (1-MA) to mature eggs for 30-60 min. Then, freshly collected sperm from male sea stars were added (1:1000 dilution) and gently mixed in the seawater. After successful fertilization, visually confirmed by the presence of a fertilization envelope, the glass dishes containing growing embryos were placed in incubators maintained at a temperature of C. Water changes () in the glass dishes were carried out weekly, removing the buildup of waste.
Sea star larvae were fed with red algae (Rhodomonas lens) and green algae (Dunaliella tertiolecta), obtained from the National Center for Marine Algae and Microbiota (NCMA, Bigelow, ME). Algae were cultured in 500 mL Erlenmeyer flasks covered with aluminum foil containing filtered seawater (FSW) and nutrients (Micro Algae Grow, Florida Aqua Farms) in 1:100 dilution of nutrient media in FSW, and kept in an incubator maintained at a temperature of C, under a 10:14 hour dark:light cycle using LED lighting. Algae flasks were shaken at least once daily to prevent settling at the bottom. The algae from these flasks were fed to the larvae in 1:500 concentrations (algae:seawater). The algae cultures were split weekly by diluting the algae culture into 4 parts by volume of fresh seawater and nutrients.
The sea urchin larvae (Lytechinus variegatus) 4-5 days post-fertilization were obtained in 100 mL falcon tubes from the animal culture facility at The National Resource for Aplysia on the campus of University of Miami’s Rosenstiel School of Marine, Atmospheric, and Earth Science. The larvae were transferred to 1 L beakers with seawater, and were kept in the lab for 1-2 days during experimentation. They were fed red algae (R. lens) and green algae (D. tertiolecta), and the beakers were maintained at room temperature (C), with constant lighting.
@@
For experiments, a randomly chosen larva was confined or squeezed gently between a bottom glass slide and a top cover slip with different heights (z-direction) using double-sided spacer tape of precise thickness (Nitto Inc., PET-based Double-coated Tape No. 5605, 5615, UTD-10B), and fluid flow-fields were quantified in a quasi-2D x-y plane. Slides were prepared by attaching a pair of long rectangular strips of double-sided tape of different thicknesses 50/100/150 m on the slide edges to serve as spacers for a required boundary height () (Figure 1). The tested boundary height () values ranged from 50-850 (Supplementary Table S1).
A 50 L sample of seawater containing the larva was transferred using a P200 micropipette from the culture dish into the center of the prepared glass slide with the spacer tape walls at the edges. For flow visualization, 10 of a 1:100 dilution of tracer polystyrene microspheres (Polysciences Inc, cat. no. 07310-15) in seawater was added to the droplet with the larva and gently mixed.
A cover slip was slowly introduced onto the glass slide containing the larva and microspheres and attached to the spacer tape walls. Live-imaging experiments were carried out by mounting this slide on an upright microscope (Zeiss Axio Imager M2). The microsphere movements were imaged using darkfield time-lapse imaging using a 10X objective to capture a 1.5 mm 1.5 mm field of view. The time-lapse images were acquired at 10 fps for a duration of 30 s using a high-speed camera (Hamamatsu ORCA-Fusion Gen-III sCMOS) attached to the Zeiss microscope. Three recordings of each organisms were carried out. The best quality videos (i.e. no bead clumping, minimum larval displacement, no larval contraction, and larvae located in the middle of field-of-view away from the walls) were selected for analysis (Supplementary Table S1).
@@
The larval flow fields were qualitatively visualized using flowtrace technique in Image-J [32]. The analysis was run for 30 s durations over a time-averaging window of 3 s. Flowtrace visualizations were carried out on all the experimental datasets, and these videos were used to count the number of vortices (number of closed circular pathlines) for each (Table S1, Supplementary Information). The number of vortices was noted for the entire 30 s time sequence of a single dataset. In instances where there was high variation in the total number of vortices over 30s, the average number of vortices was calculated as a weighted average over the total number of frames. Else, the average number of vortices was assumed to be the mode. Weighted average for number of vortices was used primarily with the early star data due to larval contractions and displacements. The data were aggregated by organism for later statistical analysis (Table S1, Supplementary Information).
The Particle Image Velocimetry (PIV) technique was adopted for quantification of the experimental larval flows. The time-lapse images were post-processed using PIVlab 2.62 in MATLAB R2020b [33]. The time series were first imported into PIVlab’s GUI via the ’time resolved’ Image sequencing style. Image pre-processing was done by enabling the Contrast Limited Adaptive Histogram Equalization (CLAHE) of 64 pixels with auto contrast stretch. A mask was created to exclude regions where a larva was present, ensuring that only the flow region was analyzed. A Fast Fourier Transform (FFT) based cross-correlation algorithm was used to compute the velocity fields. For the interrogation window settings, a multi-pass cross-correlation algorithm was employed with an initial window size of 100×100 pixels, and a second pass of 50×50 pixels, with 50 overlap between windows to improve spatial resolution. After the analysis, the datasets were converted to physical units by applying a calibration factor based on a known reference scale in the image. Post-processing of the vector fields was carried out by validation the vectors using a standard deviation threshold of 8, and a local median filter threshold of 3. Missing or invalid vectors were interpolated using local median filtering. Additionally, a velocity threshold was applied to discard unrealistic velocities exceeding the expected range based on the experiments. The final images were generated by applying ’smooth data’ filter to the highest level and high pass vector field strength to the lowest level. The same settings were applied for all the image sequences and mean was calculated for a 3 s time window. The entire analysis was repeated for multiple datasets to ensure consistency and reliability. Flow parameters such as vorticity or velocity magnitude were plotted on a 2-D, x-y plane with velocity vector fields. The average vorticity and velocity magnitude were calculated over a time window of 3 s to obtain smooth fluid flow visualizations and quantification. Finally, the distance from the center of vortices to the larval body surface was also measured (Figure S2, Supplementary Information).
Statistical Analysis
In instances where multiple recordings occurred over one individual, dependent variables were averaged with respect to the individual to ensure independent measurements. This assumed that recordings taken from the same individual were comparable. Dependent variables analyzed included number of vortices, distance from organism to vortex center (), average vortex circulation, average vortex size, average vorticity, and average vortex flow speed.
Data was compared with respect to two factors – body plan (early stage sea star, late stage sea star or sea urchin), and confinement (). To observe initial trends with respect to confinement (), data were linearly regressed with respect to .
For more detailed statements on the relationship between flow fields, body plan, and confinement, confinement – measured on a continuous scale – was divided into three strengths: strong (0 H/c ), moderate ( H/c ), and weak ( H/c 1). Aggregating points allowed for organisms with different flow field patterns (i.e., number of vortices due to morphology such as early stage sea star and sea urchin) to be compared at similar confinement strengths.
Data were not normally distributed and often violated the assumption of equal variance between factors (body plan and confinement strength) required for parametric analysis, and thus were analyzed using nonparametric means. Nonparametric Kruskal-Wallis tests and Steel-Dwass All-Pairs were used to analyze data with respect to body plan, confinement strength, confinement strength within body plan, and body plan within confinement strength. Statistical analysis was performed in JMP Pro 17.0 (SAS Institute Inc., Cary, NC).
@@
To theoretically model fluid flows generated by the ciliated marine larvae under squeeze-confinement, we used an approach utilizing superposition of Stokeslets [7, 6]. In the experiments, ciliated larvae were confined between two plates and the flows were measured at the midplane, i.e. z = . Recently, flow fields generated by a Chlamydomonas confined between two parallel plates was studied using both experiments and theory [7]. Inspired by this work, we extended the theoretical model and applied it to our experiments in ciliated marine larvae. The theoretical model starts with the incompressible 3D Stokes equation (as in Ref [7]),
@@
where is the point source of force generation in the z = plane. The solution to the above equation (using Fourier analysis) is obtained as:
The velocity fields are numerically solved over a 30 x 30 x-y grid in MATLAB to obtain the theoretical flow fields. In Ref [7], the Chlamydomonas has two flagella that beat to generate vortices, however, here the marine larvae generate fluid vortices using a large number of cilia located on their ciliary bands. Although the methods of inducing fluid flow are different in the two studies, the consequence, i.e. vortex generation, is the same in both cases. Here, the focus is on modeling the overall flow fields induced by confinement rather than the local mechanism of vorticity generation.
@@
PDF file with supporting text, Figures S1–S6, Tables S1–S12, and Video Legends.
Videos S1 and S2.